OpenStudy (anonymous):

Use part I of the Fundamental Theorem of Calculus to find the derivative of integral from x to 5 of sin(t^2)dt I will also show the equation written out below

OpenStudy (anonymous):

Use part I of the Fundamental Theorem of Calculus to find the derivative of the integral \[F(x)=\int\limits\limits_{x}^{5}\sin(t^2)dt\]

OpenStudy (anonymous):

\[\int\limits_{x}^{5}\sin(t^2)dt=-\int\limits_{5}^{x}\sin(t^2)dt\]

OpenStudy (anonymous):

so F'(x)=-sin(t^2)

OpenStudy (anonymous):

so in terms of x would you just write -sin(x^2)?

OpenStudy (anonymous):

so F'(x)=-sin(x^2)

OpenStudy (anonymous):

yes, my mistake

OpenStudy (anonymous):

Thank you so much! I appreciate the help

OpenStudy (anonymous):

yw